OpenStudy (anonymous):

Medal!!!! and Fan!!! PLEASE help!!!! 1.Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4 A.y – 7 = 4(x – 3) B.y – 7 = 4(x + 3) C.y + 7 = 4(x + 3) D.y + 7 = 4x – 12 2.Which is an equation in point-slope form for the given point and slope? Point: (2, -6); Slope: 4 A.y-6= -3/4 (x+2) B.y-6= -3/4 (x-2) C.y+6= -3/4 (x+2) D.y+6= -3/4 (x-2)

OpenStudy (cwrw238):

the point slope form is y - y1 = m(x - x1) where m = slope and (x1,y1) is the given point

OpenStudy (cwrw238):

for question 1 replace m by 4 y1 by 7 and x1 by -3

OpenStudy (cwrw238):

question 2 is similar

OpenStudy (anonymous):

so it would be y-7=4 (x-3)

OpenStudy (cwrw238):

no be careful of the signs x = -3 so the parentheses part is (x - (-3))

OpenStudy (anonymous):

Oh okay thank you! and 2 would be..... hold on im gonna try to figure out 2 and tell me if im right, Thank you so much im starting to get it :D

OpenStudy (cwrw238):

theres something wrong with question 2:- if the slope is 4 then -3/4 is wrong

OpenStudy (cwrw238):

maybe the slope should be -3/4 ( in the question that is)

OpenStudy (cwrw238):

B is the correct answer to question 1

OpenStudy (anonymous):

i can't figure out question 2 can you walk me through it please and thank you :)

OpenStudy (anonymous):

i think 2 is a but im not sure

OpenStudy (cwrw238):

as i say the answers don't match the question if we assume that the question is wrong (about the slope) and options are right then we replace m (the slope) by -3/4, x1 by 2 and y1 by -6 we get y -(-6) = -3/4(x - 2) y + 6 = -3/4 (x - 2)

OpenStudy (anonymous):

OHH sorry the slope for number 2 is -3/4

OpenStudy (anonymous):

so if the slope is -3/4 th answer would be A.?